Sodankylä
Geophysical
Observatory
Oulun yliopisto

EISCAT Measurement Campaign 2005

Last update: 2015-8-28, 08:28:55 UTC.

2005-10-13: We packed yesterday evening already, because the night was prdicted to be hopeless. The data backup is in its final stage and soon we'll be heading home!


2005-10-12: Raining all day, and discovered a new weather phenomenon: clouds so low that you could in fact see the top of the mountain, but not the bottom. In the evening clear in Kiruna, so DAEA was run on the UHF.


2005-10-11: This morning it was raining like mad with very strong winds; we thought it might blow the entire Hilton away. The mountain was not visible today, so I guess the weather can be considered "not good." Also the forecast for tomorrow looks bad.


2005-10-10: This morning it's cloudy, but no rain anymore. Actually, the clouds are kind of bright-grey, which means the weather's actually really nice! We can see the mountains, and the highest is covered by snow now, finally it seems to get colder. In the evening rain again, but good conditions in Kiruna, so time to run DAEA.


2005-10-09: The day starts beautiful, but then ends in rain and more rain. However, last night we finally got what we wanted: there were several beautiful displays of the aurora during a very clear night in Tromsø. Of course it hits us now, because the Finnish part of the campaign is over. However, the UK experiments were running fine and our optical instruments are still deployed, so we got a lot of good, joint data.


2005-10-08: Today is a beautiful day here, blue sky and the sun shining. A day to lift our spirits after the long cloudiness and rain. Also a good day for driving: Markku and Antti, Mikko, Ilkka, and Timo, as well as Kari leave for Sodankylä and Oulu. Tero and Thomas continue the optical operations until Thursday.

Last night was initially bad, but then it turned to the better. There were bright aurorae on the sky at midnight local time (22 UT), noticable through the clouds. Kilpis was clear though. The German HFSEE experiment had the higher priority tonight, so we had to wait until 22 UT to take over the UHF radar. We ran arc1. During the experiment, i.e. from 2330UT the sky was clear in Tromsø. The main auroral activity was in the south, but arcs came back north between 0030UT and 0300UT. We missed the built-up to this event, but we certainly caught a significant part of it. (Photo: Northern Lights over EISCAT)

Optical: The photometer has problems to get timing signal form the GPS receiver since the evening of 7th Oct. The AP7 set up was 2 s exposure plus 8 s exposure delay plus download time. The TV camera record started around midnight. We saved it for active moments because after about an hour of operation the picture goes bad.

Tomography: At the moment there are 4 active Russian satellites, which are on Planes 102, 104, and 106. Plane 104 includes two different satellites whose inclination differs a little.

The stations of the Finnish Tomography Chain are located at their usual locations (Kilpisjärvi, Kiruna, Luleå, Kokkola, and Nurmijärvi) during this campaign.

In collaboration with Mike Rietveld and Brett Isham, staff of PGI (Murmansk) runs a Russian tomography measurement campaign related to EISCAT heating experiments and scintillation effects parallel to the Finnish-UK EISCAT campaign (Photo: Russian receiver). They are operating three stations: one at the Tromsø EISCAT site, a second station in the south and a third in the north of the radar site. Moreover, the Radio and Space Physics Group of University of Wales Aberystwyth has their ionospheric tomography receiver at the EISCAT site (Photo: Welsh receiver).


2005-10-07: You'll never guess: it wasn't raining this morning! It's cloudy though, but there are some distinctly brighter patches on the sky and visibility is good (we can see the mountains). It was a busy day due to Antti's heating experiments, which were split into two blocks. They were interesting to run and successful, esp. the second one was not boring at all since we needed to hunt the conducting layer with the remote radars by changing the common-volume. In the evening there was a six-hour run for HFSEE, so we started our experiment only at 2200 UT. There were plenty of aurorae on the sky, visible through the clouds. Clear above Kilpis, so we might have a very successful last night.

Radar operations: We are running ANTII part 1 now, using UHF and VHF radar together with the heater. We started with arc_dlayer and arc_dlayer_ht and then switched to sippiD12. There were a number of crowbars again, the UHF is still unstable. Later in the afternoon we ran another set of heating experiments: first ten minutes of beam swinging and then another heating modulation for the benefit of the VLF group. Apparently, Tauno and Jyrki celebrated every new spectral line they could hear from the heater, which must've been quite some party. In parallel, we ran sippiD12 on the VHF and CP1L on the UHF with the remote sites looking first at the E-layer peak, later at the F2-layer peak. In the evening we ran arc1 on the UHF, starting at 2200 UT after the HFSEE, which had the higher priority. The Cluster run had the lowest priority and was not run.


2005-10-06: What do you think the weather's like this morning? It's raining of course! The evening started with holes in the cloud cover and auroral activity; by now it's cloudy again, and the E layer has vanished according to the UHF radar.

Radar operations: ANTII is cancelled for today, i.e. no VHF and Heater run 0930UT to 1130UT. We hope that the problems with the UHF radar have been resolved at least in part; we can run at just under half power (1070 kW according to RTG). After a number of crowbars at the beginning the experiment runs stably now.

VLF: Jyrki sent an e-mail this morning with some data of 4th October. Another beautiful example how well the new VLF receiver works.


2005-10-05: Well, guess what? It's grey and raining this morning. At the end of the working day (15.15LT), it's still raining.

Radar operations: ANTII is running for the first time with the VHF radar. After some initial troubles we are up and running since about 0935UT. The heating cycle started at 0950UT and alternates between two 20-min cycles of short and long heating periods. The VHF initially ran a version of arc_dlayer, which is modified by Ilkka to suit the frequencies of Markku's/Jussi's receiver. Later on, we switched to sippiD12. Due to favourable conditions, we decided to continue for another set of two hours VHF+Heating, interrupted by some 45 min of testing of the VHF by the EISCAT staff.

Brett Isham and Mike Rietveld intended to run their campaign with VHF, UHF, and possibly the heater from 1600UT to 2200UT and we were supposed to take over any time after 2000UT. However, for some reasons both radars refused to play and we had to give up on the idea. This is good for us, because we save the radar time from an otherwise completely rainy night. However, it might be bad for us in the long run, if the problems turn out to be serious.

Optical: Scanning and zenith photometers have worked well since the 630 nm channel has been disconnected. The +5 V voltage is taken from the power supply of the zenith photometer. Due to bad weather, a re-wiring has not yet been attempted.

The digital AP7 camera has been running with 30 s sampling and 2 s exposure; computer and camera need to be rebooted to get the camera working in the past two days.

The TV camera is still running with the spare lens (17 mm, f/3.5). The new 7.2 V rechargable battery (Sanyo N-6PT) has arrived, but it is not yet in use, we are waiting for suitable weather for testing. After about an hour of operation, the picture appears "frozen;" it shows a memory effect of the image amplifier. Just the brightest lights can be seen through the frozen picture.


2005-10-04: It rained all night, partly very heavily, and it's still raining now, mountain tops hardly visible in Tromsø. Good news (maybe) for optical operations: the rechargable battery for the auroral TV camera arrived today. Björn called at about 2000 UT reporting a very faint arc overhead in Lyngen. By 2030 UT what was a beautiful starry night had changed to cloud cover. Also Kari reports more and more haze and clouds in Kilpisjärvi after a beautiful sunny day. Good news from the photometer: Tero managed to fix it somewhat. The problem seems to have been in the 630-nm channel, which is now disconnected. In the course of the night, it became cloudy at all sites, and with the clouds there came more rain.

Radar operations: Time to run ANTII: the VHF is still not available, and the UHF shows the same problem as last night, no transmission despite everything looking good. The problem was solved, something (line driver?) had burnt in the Tx electronics. Experiment finally started at 1045 UT with the radar; the heating cycle started at 1055 UT and will end at 1255 UT. The UHF will continue until 1307 UT. Several experiments are being tested. Please refer to the MasterPlan™ for details.

The evening UHF experiment will start with testing Markku's and Ilkka's experiment sippi4. For this we will switch every half hour between running sippi4 and arc1 for comparison. The first interval will be sippi4. Please refer to the MasterPlan for details of experiment durations. The radar worked fine until about 2310UT, then a power failure in the Tx forced us to stop for the night.

Radar cancellations: The following experiments are cancelled:
- 3.10.-4.10. 1830-0130UT cancel both ESR and VHF
- 5.10.-6.10. 1830-0130UT cancel both ESR and VHF
- 6.10.-7.10. 1830-0130UT cancel both ESR and VHF
- 7.10.-8.10. 1830-0130UT cancel both ESR and VHF


2005-10-03: Apart for a hole in the clouds around 1800 UT, it was cloudy and raining all night. The photometer does not work, the digital auroral camera (AP7) does not work due to overheating. The booster, which gets very warm during operation, was temporarily removed and the camera left switched off for cooling down.

Radar: Transmitter (UHF) and experiment console (Eros) did not indicate any error, everything looked just fine, but for some reason no pulses were transmitted. After trying to restart Eros, reboot the crate computer, and switching various standard experiments, we decided to give up and shut down the radar after some hour and a half.


2005-10-02: Tonight we will run the UHF as planned, using arc1. Markku, Mikko, and Ilkka tested their new Sippi4 experiment during an extra time slot of 1815 to 1828 UT. Since 1830 UT arc1 is running, Mikko and Ilkka will take care of this night's run. In parallel, raw data is being recorded by Jussi's receiver system. Clear in Kilpis, cloudy and raining in Tromsø. Photometer broken again. Kari reports a diffuse arc in the north of Kilpis, getting brighter (2025 UT); the radar sees it soon after. Very foggy in Tromsø again, night ends without further complications.

Radar: Due to VHF not being available and due to Mike being busy on Monday, we decided to cancel the ANTII experiment on Monday (0930-1130 UT). Monday evening we will run the UHF as scheduled (1830-0130+UT; confirmed).

VLF: Morning chorus was very strong and lasted well beyond noon. After 1700 UT whistler activity started. It had a maximum around 1830 UT when more than 20 whistlers per minute were recorded. Tauno drove to Sodankylä and will return on Monday. At 2053 UT auroral hiss. Whistler activity is still high (>20 per minute). No observations of optical aurora, it is raining. At 2145 UT chorus emissions related to pulsating aurora started. Whistler activity high.


2005-10-01: Today no radar experiments. Kari visits Tromsø and repairs the photometer. Markku, Mikko, and Ilkka investigate Jussi's receiver setup and find that they can use it for their experiments.

VLF: Although the night was quiet some dawn chorus appeared. Daytime was quiet, but after 1900 UT we observed first auroral hiss, and after 2200 UT chorus emissions related to pulsating aurora started. They had very peculiar features which require detailed study. Some whistlers were also observed. Most of the day the computer was analysing the data and making colour plots.


2005-09-30: In the morning it was cloudy, now half-cloudy. Tero, Antti and Ilkka are on an expedition to the mountain behind the radar. Timo is on the way to Svalbard. We did not manage to get a replacement battery for the TV camera in Tromsø, but we ordered one from Sweden, which will hopefully arrive on Monday, but maybe only on Tuesday. The long-range weather forecast is a bit hopeless, rain for the next days, i.e. optical-campaign weather. ;-( Yesterday's halo-CME was a false alert, it blasted off the backside of the Sun away from Earth. Bad news from the Tromsø photometer: the counters only see zeros, something's obviously wrong.

Radar operations: It seems the VHF will work, but only on one half of the antenna. Also, the VHF might need to borrow a card from the UHF to work, which means that the UHF won't work. This is BAD, because the idea was to run VHF and UHF simultaneously. EISCAT staff is still working on the problem. Schedule changes: ANTII is shortened on every day (3.-7.10.) from originally 0830-1230UT to 0930-1130UT.

VLF: Nice dawn chorus lasted until 09 UT. We had an excursion group from SGO visiting. They were very interesred in our receiver system, and it seems that they also were satisfied with the visit (especially due to the chance to shoot with Tauno's rifle). The rest of the day was very quiet except that we learned a lot of sferics and Earth-ionosphere waveguide modes and cutoffs. With the new receiver we have the possibility to detect sferics as they are instead of them being overflows in the amplifiers.


2005-09-29: Sunny in Tromsø. Last night it cleared up, but then almost immediately temperature fell and resulted in dense fog. TV camera is causing still problems due to worn out rechargable battery. We'll try to get a replacement today. Kari reports photometer correctly aligned and TV camera working, digital camera will be put into place next night.

Radar operations: VHF will most likely work! This information together with discussions with Kari and Antti lead to the following schedule changes: ANTII experiment cancelled on Sat and Sun (1.10., 2.10.; 0830-1230UT) due to other operations at Heating. UHF-only experiment cancelled on Sat night (1830-0130UT) due to bad weather and low(ish) activity predicted. Three-radar experiments confirmed on Fri night and Sun night (30.9., 2.10.; 1830-0130UT; UHF+VHF+ESR). See MasterPlan.

Mikko, Ilkka, Timo, and Antti arrived alright in Tromsø. Suprisingly, clear skies still prevail. A small halo-CME was detected earlier today.

VLF: Morning chorus started about 0300 UT and lasted until noon. Evening was quiet. The first plots from this campaign were made, which was easy due to digital recording.


2005-09-28: Sunny in Tromsø, but it does not last very long. UHF and VHF down today, staff tries to get both radars ready for our campaign. Kari leaves for Kilpis.

Telephone discussion with Antti: we will run the day-time heating experiments (ANTII) even if VHF is not available. We'll try to substitute it by UHF, if that fails then only dynasonde (API). Antti prefers 4x2 hrs instead of 2x4 hrs (6x2 hrs if w/o radar), so the scheduled times are likely to be changed from 0830-1230 UT to 0930-1130 UT.

VLF observations: Dawn chorus continued right away after pulsating aurora. It was occasionally very strong and lasted until noon. In the afternoon about half an hour 'silence,' but then chorus appeared again. Evening was quiet except some dircrete emissions occurred occasionally.


2005-09-27: Campaign weather in Tromsø: raining and low visibility. Optical instruments installed in the afternoon. All is well, but the TV camera seems to be damaged. Attempting some workaround. Some aurora visible between the clouds.

VLF observations: Moderate dawn chorus until 16 UT, then no activity until 22 UT when chorus related to pulsating aurora started. Around midnight several simultaneous very strong chorus emission bands with 'PLHR control.' Many new features were observed.

2005-09-26: Kari, Tero, and Thomas travelled to Tromsø via Karesuvanto and Kilpisjärvi, where optical instruments were set up. Due to daytime and weather conditions, exact pointing of equipment will be fixed later. Tero installed tomography antenna at KIL, station tested and working.

Jyrki and Tauno hunting (VLF waves and birds) in the Future (642). VLF recordings started at 1231 UT. Some events were observed.


Last modified: 28 August 2015, 11:28:55.